How do I write the equation of the line that passes through point (-2,3) with a slope of -4?
Answer by broadway(2)

we are trying to get the equation of a line which is y=slope(m)x+y intercept(b). so, y=mx+b. we have y, m, and x. so, just plug these in and solve for b!
3=-4(-2)+b
3=8+b
-5=b
so, your equation is y=-4x+-5